OpenStudy (anonymous):

On the graph of the equation 5x – 2y = 20, what is the value of the y-intercept?

OpenStudy (anonymous):

\[5-2dy/dx=0\] \[dy/dx=5/2\]

OpenStudy (anonymous):

-10 is your y intercept

OpenStudy (anonymous):

y=5/2x-10

OpenStudy (anonymous):

Rewrite the equation in the form y = mx + b, where b is the y-intercept. -2y = -5x + 20; y = 5x/2 - 10; so, b = -10. Your y-intercept is therefore (0, -10).